If Frank has 85 chickens in stock and he forecasts the following monthly demands. Using a level production plan how many chickens will Frank need to supply each month to meet the demand?
January = 190
February =272
March = 170
April = 361

Calculate your answer to one decimal place.
Answer #1

The net total requirement for the four months = (190+272+170+361) - 85 = 908

The production level = 908 / 4 = 227

Month Demand Production Ending Inventory
85
Jan 190 227 122
Feb 272 227 77
Mar 170 227 134
Apr 361 227 0

So,

Frank need to supply each month 227 chickens to meet the demand.